OpenStudy (anonymous):

I got an answer but i need help for interval. A particle moves along a straight line and its position at time t is given by s(t)=2t^3−30t^2+126

OpenStudy (anonymous):

Use interval notation to indicate the time interval or union of time intervals when the particle is moving forward and backward. Forward: backward Use interval notation to indicate the time interval(s) when the particle is speeding up and slowing down. Speeding up: slowing down:

OpenStudy (anonymous):

\[s(x)`=6(t-3)(t-7)\] and s(t)``=\[12(t-5)\]

OpenStudy (anonymous):

foward i got [0,3)(7,infiity) was right

OpenStudy (anonymous):

@tkhunny please help?

OpenStudy (tkhunny):

Did you present the right equation? That factored first derivative has very little to do with the s(t) provided.

OpenStudy (anonymous):

the equation was \[2t^3-30t^2+126t\] \[t \ge0\]

OpenStudy (anonymous):

sorry about that

OpenStudy (tkhunny):

Okay, let's first talk about s"(t) - the second derivative It's linear. You should be able to determine when this is negative and when this is positive. This should be simple. It better be! Positive is speeding up Negative is slowing down.

OpenStudy (anonymous):

Oh I didnt know

OpenStudy (anonymous):

so far I understand

OpenStudy (tkhunny):

For t > 5, s"(t) > 0 and we have positive acceleration (speeding up). For t < 5, s"(t) < 0 and we have negative acceleration (slowing down).

OpenStudy (anonymous):

ok, t>5 is counted as interval?

OpenStudy (tkhunny):

\((5,\infty)\) For t = 5, nothing is going on with acceleration. s"(t) = 0

OpenStudy (anonymous):

( 5 infy) for speeding up?

OpenStudy (tkhunny):

I already said that a couple of times. Why do you continue to doubt?

OpenStudy (anonymous):

I entered it and it says it was wrong :/

OpenStudy (tkhunny):

That's a good answer. Maybe it wants \([5,\infty)\) and it considers s"(t) = 0 as some how non-negative acceleration. I really don't like online things like this, We can talk and understand all we want, but figuring out what input is required turns out to be the hard part. Very annoying!

OpenStudy (anonymous):

Yes , it is aboustely annoying. i tired [5,inf) but was also wrong :/

OpenStudy (anonymous):

any ideas?

OpenStudy (tkhunny):

Nope. I can only get the right answer. We'll have to get the test designer to tell us what it wants. It could be wrong. Maybe we're still using the wrong equation. Maybe it wants the whole thing at the same time? Slowing down: \((0,5)\) Speeding up: \([5,\infty)\)

OpenStudy (anonymous):

The equation is right, but i entered it but i keep getting wrong

OpenStudy (tkhunny):

Good luck with that. Wave a magic wand at it, I guess. Did you check the air in the tires? Really, I am sorry it isn't working. It's hard to learn math with silly stuff in your way.

OpenStudy (anonymous):

its alright, thanks for your time.

OpenStudy (tkhunny):

Maybe it wants the Forward/Backward intervals first?

OpenStudy (anonymous):

I got the answer for the forward one but not backward.

OpenStudy (tkhunny):

s'(t) = 6(t−3)(t−7) Forward: \([0,3) \cup (7,\infty)\) Backward: \((3,7)\) Again, it's not perfectly obvious how it wants us to treat t = 3 and t = 7.

OpenStudy (anonymous):

Thank you! the backward was right, as you said

OpenStudy (anonymous):

the problem is just slowing down and speeding up

OpenStudy (tkhunny):

Oh, well, there goes that theory. Punt!

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

are you familiar withLogarithmic Derivatives